OpenStudy (mica):

need help im stuck i need to prove that (1-tan x)(1+tan x) = (1-2sin^2x)(sec^2x)

OpenStudy (mica):

i forgot to mention that i have to start on the left side x.x

OpenStudy (mica):

i original problem is prove: cos2x = 1 - tan^2x ---------- 1+tan^2x

OpenStudy (jiteshmeghwal9):

\[(1-\frac{sinx}{cosx}){(1+\frac{sinx}{cosx})}\]\[(1-\frac{\sin^2x}{\cos^2x})\]\[1-\sin^2x \times \frac{1}{\cos^2x}\]\[(1-\sin^2x)(\sec^2x).\]

OpenStudy (mica):

aaaaaaaaaa lol gotcha xD thank you

OpenStudy (jiteshmeghwal9):

yw :)

myininaya (myininaya):

?

myininaya (myininaya):

\[(1+\tan(x))(1-\tan(x))\] \[1-\tan^2(x)\] \[1-\frac{\sin^2(x)}{\cos^2(x)}\] He is right until this point.

OpenStudy (jiteshmeghwal9):

? ------> ??

OpenStudy (anonymous):

hi

myininaya (myininaya):

Hint 1: Combine the fractions

myininaya (myininaya):

Hint 2: Rewrite the numerator in terms of sin by using the Pythagorean identity, \[\cos^2(x)=1-\sin^2(x)\]

OpenStudy (jiteshmeghwal9):